# Mamba-ND: Selective State Space Modeling for Multi-Dimensional Data

### 08 Feb 2024 v1

@@ 5.6 Video Specific Ablations
When initializing new SSM layers in T+T- direction, we can adjust the scale of $\Delta$ to control the initial temporal receptive field. We show such results in Table 7. We find that 1.0 is the optimal value, which coincides with the prior conclusion from S4ND [33].
@@
Table 7: Ablation Study on the Scaling Factor of $\Delta$ for HMDB51 Experiments. We report the performance on HMDB-51 under different initializations of temporal layers (T+ and T-). $\Delta=1.0$ is the optimal choice.
@@
6 Discussion
6.1 Effective Receptive Field (ERF)
To validate the significance of multi-directional design, we visualize the effective receptive field (ERF) [30] of Mamba-ND  in Figure 7. ERF is computed by setting the gradient of the center patch to one and backpropagating through the network. In particular, we visualize the ERF of ImageNet-1K pretrained weights of Mamba-ND  against the Bi-directional and Uni-directional baseline (Table 4, row 5 and Table 4, row 4). Sensitivity is measured by the whiteness of a pixel in the visualization. The Uni-Directional model exhibits a sharp cutoff and is insensitive to all patches after it in the flattened sequence. This is clearly undesirable in data where causal relations do not exist. The Bi-Directional model exhibits a global receptive field, but it is heavily biased in the horizontal direction. This is also undesirable because images can have vertical structures. The Multi-Directional model exhibits a more uniform pattern in the sensitivity visualization, which explains their superior performance.
6.2 Depths versus Widths
In Section 5.4, we showed that a complicated multi-directional design may not necessarily lead to improved performance over the alternating-directional baseline. We hypothesize that this result is related to the effective depth of the model.
7 Conclusion
Transformers have been the go-to choice in image, language, and video tasks in recent years, just like ConvNets and RNNs in earlier years. Mamba [31] presents itself as a competitive challenger to Transformers in 1D sequence modeling. However, due to their intricate formulations discussed in Section 4, it is not straightforward whether they will be equally competitive in multi-dimensional data. In this work, we proposed Mamba-ND  which successfully extends the strong performance of Mamba to multi-dimensional space. It was able to outperform Transformers on a variety of tasks with significantly fewer parameters, while reducing the quadratic complexity to a linear one. Through extensive experiments with alternative designs, we demonstrate the importance of multi-directional design in dealing with multi-directional data and highlight its advantages over uni-directional and bi-directional baselines. We hope these results will facilitate future research in multi-dimensional modeling.
8 Acknowledgement
We are grateful to Microsoft Research for supporting this research through their Accelerate Foundation Models Research program.

### 12 Mar 2024 v2

@@ 9.3 $\Delta$ Initialization
When initializing new SSM layers in T+T- direction, we can adjust the scale of $\Delta$ to control the initial temporal receptive field. We show such results in Tab. 8. We find that 1.0 is the optimal value, which coincides with the prior conclusion from S4ND [37].
9.4 Weight Inflation
There are various alternatives for inflating the positional embedding from 2D to 3D. We considered two policies. Given a 3D sequence of shape $H\times W\times T$, the first option copies the original 2D embedding of shape $H\times W$ by $T$ times and scales the values to $\frac{1}{T}$. The second option copies the 2D embedding once and places it at location $\frac{T}{2}$ on the temporal axis. All other embeddings are initialized to 0. We find no significant differences in performance between these two designs.
@@ 9.4 Weight Inflation
Table 8: Ablation Study on the Scaling Factor of $\Delta$ for HMDB-51 Experiments. We report the performance on HMDB-51 under different initializations of temporal layers (T+ and T-). $\Delta=1.0$ is the optimal choice.
@@
9.5 Ablation on 3D Segmentation
Because we train 3D segmentation models from scratch instead of from 2D-pretrained initializations, it may offer better insight into the design choices in the 3D space. However, it may also be the case that segmentation is local in nature and may require less global information. For completeness, we compare the performance of different SSM designs in Tab. 9. Results show that our simple design of alternating the scan direction between layers is still the best performing one. However, 1D-SSM with no bidirectional design also exhibits strong performance.
Table 9: Ablation Study on Layer Designs. We report top-1 accuracy on the ImageNet-1K validation set and HMDB-51 split 1. We report Dice score on BTCV dataset. The Alt-Directional design is the top-performing one.
[…]
Alt-Directional Block Level 79.4 59.0 81.5
Multi-Head Layer-Level 77.6 51.5 81.4
ND-SSM Layer-Level 77.2 46.7 80.2
1D-SSM - 76.4 34.9 80.7
Bi-SSM Layer-Level 74.6 32.1 77.9
9.6 Patch Size
One crucial element of Mamba-ND’s success depends on its capability to achieve a global context at linear complexity. This allows it to use a lower patch size or higher resolution input given fixed computation capacity when compared with self attention based methods. We highlight this in figure Fig. 7. On ImageNet-1K, we explore the effect of different patch sizes. We report the results in Tab. 10
Table 10: Ablation Study on patch size. We report the performance on ImageNet-1K. We find that lower patch size achieves better result.
[…]
ViT-B 16 77.9
DeiT-S 16 79.8
Swin-T 4 81.3
Mamba-2D-S 8 81.7
Mamba-2D-S 16 79.4
Figure 7: Mamba-ND  Computational Complexity with Respect to Sequence Length. We plot the computational cost of the standard 12-layer Vision Transformer (ViT-B), and comparable Mamba-2D with 12 and 24 Mamba layers. In terms of parameter count, 2 Mamba layers are roughly equivalent to 1 ViT layer, which includes a Multi-head Attention (MHA) module and a Feed Forward Network (FFN). $14^{2}$ reflects the standard $224\times 224$ image with a patch size of 16. Sequence length increase can either reflect a lower patch size or a higher input resolution.
9.7 Ordering
In the alternating directional setup, there is still room for exploration in the ordering of axes. For example, there is no reason to assume TWH+ is better than WTH+ for H+. There is also no intuitive reason to prefer H+ over W+ as the direction of the first layer. This means the design space incorporates $482^{3}\cdot 3!=$ 48 possible choices (there are 2 options for each of H, W, T, and there are 3! orderings of the first three layers). Due to computational constraints, we are unable to explore all options. We attempted two incremental changes from the base setup on BTCV dataset: 1) Changing the order of the first three layers to WHT and TWH. 2) Swapping the order of TWH+ to WTH+. We found no meaningful difference between the results. We hypothesize that the residual connection is sufficient to mitigate the effect of different orderings.
10 Concurrent Works
There have been recent works such as VisionMamba [56], VMamba [28], and U-Mamba [34] that focus on applying Mamba to specific application fields such as classification or segmentation. Contrary to these works, which have a limited design space, Mamba-ND  explores the generic recipe to adopt Mamba-Kernel to multi-dimensional data, which is under-explored. For example, VisionMamba can be considered as a variant of Bi-SSM with additional designs, and VMamba can be considered as hierarchical ND-SSM with additional designs.